Can you please sol e this problem and show me The steps, thank you.
Sara mixed 5 L of cranberry juice into 4 L of apple juice, that had 33% sugar. If the cranapple mixture was 40% sugar, what was the percent of sugar in the cranberry juice?

Answer by stanbon(75887)

Sara mixed 5 L of cranberry juice into 4 L of apple juice, that had 33% sugar. If the cranapple mixture was 40% sugar, what was the percent of sugar in the cranberry juice?
------
Equation:
sugarin cranberry + sugar in apple = sugar in mixture
0.33*4 + (x/100)*5 = 0.40*(4+5)
------
Multiply thru by 100 to get:
33*4 + 5x = 40*9
----
5x = 360-33*4
5x = 228
x = 45.6
---
Ans: 45.6% sugar in cranberry juice
-------------
Cheers,
Stan H.
-------------